# Two-step estimation of latent trait models

### v2

@@ 6 Conclusions and discussion
Two-step estimation is a general approach which is in principle applicable to any types of latent variable models where the focus of interest is on the structural model. In this paper we have described it for the broad family of latent trait (or item response theory) models, where continuous latent variables are measured by observed categorical indicators. In our simulations and examples, two-step estimation performed very well for them, as was also expected on theoretical grounds. We compared it with one-step estimation which estimates all model parameters at once (another possible alternative, adjusted three-step estimation, is not generally available for latent trait models). The two approaches behaved generally similarly, with two-step estimates losing next to no efficiency in large samples and often having somewhat better properties in small samples. This is in line with previous findings on two-step estimation for other families of models, in particular latent class analysis of categorical variables and structural equation modelling of continuous variables. Together these studies provide increasingly strong evidence and reassurance that the two-step method is a useful and convenient approach to estimation of latent variable models, and an attractive alternative to the existing one-step and three-step methods.
@@
We also argue that two-step estimation has two major advantages in general. One of them is practical and computational convenience, compared to one-step estimation of all parts of the model at once. This may be particularly important for latent trait models, because they are more often used with complex structural models than are latent class models, and are also inherently more computationally demanding than structural equation models. Another important characteristic of the two-step approach is that it avoids interpretational confounding, where the effective definition of the latent variables is affected by the specification of the structural model. This may be more consequential for latent class models, where the interpretation of a categorical latent class variable can change in a discontinuous way, than for latent trait models and structural equation models which have continuous latent variables.
@@
To build on these promising results so far, different directions of further research on different elements of two-step estimation could be pursued. The accessibility of the approach would be much improved by integrated implementation of it (including calculation of correct standard errors of the estimates) in general-purpose software for latent variable models, including latent trait models. This would extend and complement the procedures that are already available for latent class and structural equation modelling. For better understanding of the scope of two-step estimation, its implementation and performance could be studied for other categories of models, for example ones which combine different types of latent variables and measurement items, or for extensions such as multilevel models and ones with non-equivalent measurement models. These and other areas of further research remain to be explored.

### v3

@@ 6 Conclusions and discussion
Two-step estimation is applicable to any types of latent variable models where the measurement and structural parameters are distinct and the focus of interest is on the structural model. In this paper we have described it for the broad family of latent trait models, where continuous latent variables are measured by observed categorical indicators. In our simulations and examples, two-step estimation performed very well for them, as was also expected on theoretical grounds. We compared it with one-step estimation which estimates all model parameters at once. The two approaches behaved generally similarly, with two-step estimates losing next to no efficiency in large samples and often having somewhat better properties in small samples. This is in line with previous findings on two-step estimation for other families of models, in particular latent class analysis of categorical variables and structural equation modelling of continuous variables. Together these studies provide increasingly strong evidence and reassurance that the two-step method is a useful and convenient approach to estimation of latent variable models, and an attractive alternative to the existing one-step and three-step methods.
@@
We also argue that two-step estimation has two major advantages in general. One of them is practical and computational convenience compared to one-step estimation. This may be particularly important for latent trait models, because they are more often used with complex structural models than are latent class models, and are also inherently more computationally demanding than structural equation models. Another important characteristic of the two-step approach is that it avoids interpretational confounding, where the effective definition of the latent variables is affected by the specification of the structural model. This may be more consequential for latent class models, where the interpretation of a categorical latent class variable can change in a discontinuous way, than for latent trait models and structural equation models which have continuous latent variables.
@@
To build on these promising results so far, different directions of further research on different elements of two-step estimation could be pursued. The accessibility of the approach would be much improved by integrated implementation of it (including calculation of correct standard errors of the estimates) in general-purpose software for latent variable models, including latent trait models. This would extend and complement the procedures that are already available for latent class and structural equation modelling. For better understanding of the scope of two-step estimation, its implementation and performance could be further studied for other categories of models, for example ones which combine different types of latent variables and measurement items, for extensions such as multilevel models and ones with non-equivalent measurement models, or for other methods of estimation than likelihood-based estimation. These and other areas of further research remain to be explored.